OpenStudy (anonymous):

Find the length of the curve. y = 2x3/2 between x = 0 to x = 5/4

OpenStudy (zarkon):

use \[\int\limits_{a}^{b}\sqrt{1+(f'(x))^2}dx\]

OpenStudy (anonymous):

nice latex. did you finish? it always is cooked up so the square of the derivative is something nice

OpenStudy (anonymous):

\[f(x)=2x^{\frac{3}{2}}\] \[f'(x)=3x^{\frac{1}{2}}\] \[(f'(x)^2=9x\] so you need \[\int_0^{\frac{5}{4}} \sqrt{1+9x}\]

OpenStudy (anonymous):

ya, i did the same, square root of 1+9x . but the answer is not the same.

OpenStudy (anonymous):

antiderivative is \[\frac{2}{27}\sqrt{(1+9x)^2}\] yes?

OpenStudy (anonymous):

2/9*3*square root of 1+9x . is it true?

OpenStudy (anonymous):

think your off here.

OpenStudy (anonymous):

no,. mine is different. square root of 1+9x is at the bottom.

OpenStudy (anonymous):

you have a choice. use \[u=1+9x\] \[\frac{1}{9}du=dx\] get \[\frac{1}{9}\int u^{\frac{1}{2}} du=\frac{2}{3}\times \frac{1}{9}u^{\frac{3}{2}}\]

OpenStudy (anonymous):

you want ANTI - DERIVATIVE. looks like you are taking the derivative.

OpenStudy (anonymous):

add one to the exponent and then multiply by the reciprocal. \[\frac{1}{2}+1=\frac{3}{2}\]

OpenStudy (anonymous):

got it. thanks